Point A is at a potential of +250 V, and point B is at a potential of -150 V. An α-particle is a helium nucleus that contains two protons

and two neutrons; the neutrons are electrically neutral. An α-particle starts from rest at A and accelerates

toward B. When the α-particle arrives at B, what kinetic energy (in electron volts) does it have?


1
Expert's answer
2021-06-08T09:34:46-0400

The charge of the alpha particle is equal to the charge of two electrons (by modulus):


"q = 2|e|"

where "e = -1.6\times10^{-19}C" is the charge of electron.

By definition, the potential difference between two points "A" and "B" is equal to the work that should be done to move a unit positive charge from "A" to "B". Then the work required to move not unit charge "q" is:


"W = Vq = 2V|e|"

where "V =250V - (-150V) = 400V" is the absolute potential difference between two points.

According to the work-energy theorem, this work is equal to the change in the kinetic energy of the particle:


"K_f - K_i = W"

Since "K_i = 0" is the initial kinetic energy of the particle, the final kinetic energy is:


"K_f = W = 2V|e|"

By definition, one electronvolt is the amount of kinetic energy gained by a single electron accelerating from rest through an electric potential difference of one volt in vacuum. Thus, dividing "K_f" by the charge of the electron "|e|", obtain the kinetic energy in electronvolts:


"\dfrac{K_f}{|e|} = 2V = 2\cdot 400 = 800eV"

Answer. 800eV.